Can someone help me answer questions B and C from the last page of the lab. Please

T-Mobile 10:30 p. m. a learn.vccs.edu KINETIC AND POTENTIAL ENERGY Please at the end of the lab experiment clean your table and wait for the instructor to check you out! All the group partners must be present. Thank you. Labs goal In this experiment we will determine the Kinetic (K) and Potential Energy (U) of a cart moving down an incline. We will investigate if and how these energies change. Theory Kinetic energy is related to the motion of an object; proportional to the objects squared velocity is Potential energy is connected to the position of the object. Potential gravitational energy depends on the relative distance between the object and the center of the Earth. We use a formula that takes into account the height of the body with respect to the ground level. The sum of Potential energy +Kinetic energy is the Mechanical energy of the object. ME-U+K 1. Follow the instructors directions to prepare your table with the track, the cart, the vertical rod with the pivot clamp, the motion sensor, the angle indicator, the end stopper, the Pasco interface, and the laptop connections. 2. Incline the track at a small angle for the first trial and then increase the angle on the second trial. Choose the angles within a range of 3° - 10 3. Select short range measurement on the motion sensor
media%2Fd4a%2Fd4ac214e-f28b-412b-b590-2c
media%2Fb06%2Fb067d0ec-c49d-45c9-a719-c6
media%2F268%2F268023f8-d9dc-45bc-8ad9-6f
media%2Fa86%2Fa863fcc3-7354-4917-bfc4-85
media%2F876%2F87638370-2d46-42ce-8a86-ac
Answer #1

b)

In the experiment, Total energy is taken as sum of potential energy and kinetic energy.

There is some friction present between the cart and the inclined surface.

As the cart goes down the inclined plane, there is some energy loss due to friction which reduces that kinetic energy.

Ideally, on a frictionless surface, total energy will remain constant but here friction is present.

Hence energy loss due to friction decreases the total energy line towards the end of the cart ride.

c)

Mechanical energy plot will be different for different inclination.

As angle \theta increases, initial height h of the cart also increases, thus initial potential energy is more.

As cart starts from rest, hence initial kinetic energy is zero.

Thus, mechanical energy of the object is initial potenial energy, which changes with angle \theta .

So, as \theta increases, mechanical energy of the object increases.